### v2

# Probing quantum chaos through singular-value correlations in sparse non-Hermitian SYK model

Abstract
Exploring the spectral properties of non-Hermitian systems presents a substantial theoretical challenge due to the presence of a complex eigenvalue spectrum. Singular values for such systems are inherently real and non-negative, and the techniques used for Hermitian systems can be used with ease. As a prototypical example of such systems, we investigate the singular-value spectrum of a non-Hermitian extension of the sparse Sachdev-Ye-Kitaev (SYK) model, a solvable toy model of quantum chaos and quantum gravity with significant interest in digital quantum simulation. Our findings reveal a congruence between the statistics of singular values and those of the analogous Hermitian Gaussian ensembles. An increase in sparsity results in the model deviating from its chaotic behavior, a phenomenon precisely captured by both short and long-range correlations in the singular-value spectrum. These findings indicate the existence of a critical sparsity threshold, beyond which the chaotic nature of the model breaks down, suggesting a potential collapse of holographic duality in such systems.
YITP-24-70, RIKEN-iTHEMS-Report-24
Introduction: Spectral statistics serve as a critical tool for probing the energy levels in quantum systems, offering insights into the dynamics within quantum chaos and Random Matrix Theory (RMT) Mehta (1991); Bohigas et al. (1984); Wigner (1955, 1957); Dyson (1962a, b); Berry et al. (1977); Berry and Tabor (1976). Short-range level statistics such as the ratio of consecutive level spacings Atas et al. (2013) and the long-range correlations such as the Spectral Form Factor (SFF) Agam et al. (1995); Guhr et al. (1998); Leviandier et al. (1986); Wilkie and Brumer (1991); Brézin and Hikami (1997); Cotler et al. (2017) reveal the symmetries and randomness in the system. The latter has also been understood from the semiclassical gravity Saad et al. (2018) and measured in many-body scenarios Das et al. (2024); Dong et al. (2024). Analogously, the upsurge in research has been directed towards unraveling the quantum chaotic properties of open quantum systems Denisov et al. (2019); Yang et al. (2024); Nandy et al. (2024) described by non-Hermitian Hamiltonians Bender and Boettcher (1998); Shibata and Katsura (2019); Ashida et al. (2021); Cornelius et al. (2022); Matsoukas-Roubeas et al. (2023a). This connotes a departure from conventional methods suited for Hermitian systems and necessitates new methods for spectral analysis Sá et al. (2020); Li et al. (2021, 2024); Cipolloni and Grometto (2024); García-García et al. (2023a); Xu et al. (2021); Zhou et al. (2024). One such method approaches through the Singular Value Decomposition (SVD) Kawabata et al. (2023), which facilitates a straightforward generalization of the frameworks used for Hermitian systems with real and non-negative singular values. Consequently, this enables us to define the singular spacing ratios Kawabata et al. (2023) and singular form factor ($\upsigma$FF) Roccati et al. (2024) in a much more elegant manner capturing transitions from integrability to chaos, many-body localization Hamazaki et al. (2019); Roccati et al. (2024), and the non-Hermitian skin effect Martinez Alvarez et al. (2018); Yao and Wang (2018); Hamanaka and Kawabata (2024). Outside this domain, SVD has also found applications in generalizing entanglement entropy for non-Hermitian transition matrices in AdS/CFT correspondence Parzygnat et al. (2023).
In this paper, we utilize the SVD framework to study the quantum chaotic properties and their transitions in the sparse, non-Hermitian version of the Sachdev-Ye-Kitaev (SYK) model Sachdev and Ye (1993); Kitaev (2015). Standing as a quintessential archetype for quantum chaos and possessing a dual gravitational theory at low temperatures Maldacena and Stanford (2016), the SYK model (see See Chowdhury et al. (2022); Rosenhaus (2019) for comprehensive reviews) can be adapted into sparse variants Xu et al. (2020); Tezuka et al. (2023); Anegawa et al. (2023); Caceres et al. (2021); Cáceres et al. (2023), maintaining its chaotic nature up to a critical sparsity threshold García-García et al. (2021); Orman et al. (2024). This adaptation holds theoretical considerations in holography as well as the practical significance for the deployment of the SYK model on quantum processors Jafferis et al. (2022); Kobrin et al. (2023), where managing a sparse matrix is more viable. Our focus is particularly drawn to the influence that sparsity exerts on the average singular value spacing ratio, denoted by $\langle r_{\upsigma}\rangle$, and the $\upsigma$FF, which serves as an extension of the conventional $\braket{r}$-value Atas et al. (2013) and the SFF in Hermitian systems, respectively. Additionally, we introduce singular complexity, inspired by spectral complexity, a boundary dual to the volume of the Einstein-Rosen bridge in gravitational theories Iliesiu et al. (2022). Our findings suggest that these metrics proficiently encapsulate the transition from chaos to integrability in the non-Hermitian SYK model, as we outline in the following sections.
Non-Hermitian sparse SYK model: The prototypical model we consider is the $4$-body non-Hermitian sparse SYK model (nSYK) Sachdev and Ye (1993); Kitaev (2015), given by the following Hamiltonian
$H_{\mathrm{nSYK}}^{\mathrm{sparse}}=\sum_{1\leq a<b<c<d\leq N}x_{abcd}(J_{abcd}+iM_{abcd})\,\psi_{a}\psi_{b}\psi_{c}\psi_{d}\,.$ (1)
where $J_{abcd}$ and $M_{abcd}$ are the independent random couplings drawn from the Gaussian ensemble with zero mean and variance $\braket{J_{abcd}^{2}}=\braket{M_{abcd}^{2}}=6/N^{3}$ (see Lau et al. (2021); Ozaki and Katsura (2025) for clean SYK model). The couplings $M_{abcd}$ explicitly break the Hermiticity of the Hamiltonian. The variables $\psi_{k}$ are the Majorana fermions obeying Clifford algebra $\{\psi_{a},\psi_{b}\}=\delta_{ab}$. Additionally, the random variables $x_{ijkl}\in[0,1]$ introduce an element of sparsity to the Hamiltonian. They follow a Gaussian distribution with zero mean and a variance of $\braket{x^{2}_{abcd}}=6/(pN^{3})$ Xu et al. (2020); García-García et al. (2021). The parameter $p$ governs the probability of $x_{abcd}$ being unity, thus determining the sparsity level within the Hamiltonian. A fully dense model is characterized by $x_{abcd}=1$ or $p=1$ for all combinations of $\{a,b,c,d\}$, aligning with the original nSYK Hamiltonian García-García and Godet (2021); García-García et al. (2022a, b); Cipolloni and Kudler-Flam (2023). Conversely, a completely sparse model would have $x_{abcd}=0$ or $p=0$ for all combinations, representing the other extreme Xu et al. (2020). The addition of sparsity effectively reduces the total number of independent terms in the Hamiltonian from $O(N^{4})$ to $O(pN^{4})\sim O(kN)$, where $k$ is defined in (4). Yet, the probabilistic nature of $x_{abcd}$ implies that the exact number of non-zero terms in $H_{\mathrm{nSYK}}^{\mathrm{sparse}}$ is not fixed but rather determined by the probability $p$. We follow Orman et al. (2024) without fixing the total number of terms in the sparse Hamiltonian in each realization.
Singular value decomposition and singular value statistics: In Hermitian systems, the eigenspectrum of the Hamiltonian is a key indicator of chaotic dynamics. The eigenvalues $E_{n}$ are real, allowing for the definition of consecutive level spacings $s_{n}=E_{n+1}-E_{n}$, as an ordered list. For large Hilbert spaces, the average consecutive level spacing ratio $r_{n}$ is defined as $r_{n}=\mathrm{min}(s_{n},s_{n+1})/\mathrm{max}(s_{n},s_{n+1})$ with $\braket{r}=\mathrm{mean}(r_{n})$ being the average $r$ value Atas et al. (2013). Its usefulness lies in simplifying the numerical analysis by bypassing the need to unfold the spectrum for level statistics. While $\braket{r}$-value is well-defined for Hermitian systems, it falls short in applying for non-Hermitian systems where eigenvalues are complex, and thus an ordered sequence of level spacings is not feasible. In such scenarios, one may consider complex spacing ratios Sá et al. (2020) and correspondingly the statistics of complex eigenvalues. However, an alternate proposal suggests focusing on singular values derived from the singular value decomposition (SVD) of the Hamiltonian Kawabata et al. (2023). SVD is expressed as $H=U\Sigma V^{\dagger}$, where $U$ and $V$ are unitary matrices, and $\Sigma$ is a diagonal matrix containing the real and non-negative singular values $\sigma_{i}$ Nielsen and Chuang (2010). In Hermitian matrices, singular values are simply the absolute values of the eigenvalues, whereas, in non-Hermitian matrices, their relationship is more complex Allard and Kieburg (2024).
The singular values of a non-Hermitian matrix $H$ can be computed by the Hermitization process Feinberg and Zee (1997), which involves constructing a new matrix $\mathsf{H}$, twice the dimension of $H$. This larger matrix incorporates both $H$ and its Hermitian conjugate $H^{\dagger}$ in a block structure Feinberg and Zee (1997)
$\mathsf{H}=\begin{pmatrix}\mathbb{0}&H\\ H^{\dagger}&\mathbb{0}\end{pmatrix}\leavevmode\nobreak\ \leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ \leavevmode\nobreak\ \mathsf{H}^{2}=\begin{pmatrix}HH^{\dagger}&\mathbb{0}\\ \mathbb{0}&H^{\dagger}H\end{pmatrix}\,.$ (2)
The eigenvalues of $\mathsf{H}$ are {±$\{\pm\sigma_{n}\}$, precisely containing the singular values of $H$. This is equivalent to constructing an effective Hamiltonian $H_{\mathrm{eff}}=\sqrt{H^{\dagger}H}$ or $H_{\mathrm{eff}}=\sqrt{HH^{\dagger}}$ so that the eigenvalues of $H_{\mathrm{eff}}$ represent the singular values of $H$ Feinberg and Zee (1997). Such construction of a Hermitized matrix resembles the formulation of the Wishart SYK model and its supersymmetric version, comprised of non-Hermitian conserved charges or supercharges Li et al. (2017); Sá and García-García (2022).
Given the singular values $\{\sigma_{i}\}$ of the corresponding non-Hermitian Hamiltonian $H$, one analogously defines $\lambda_{n}=\sigma_{n+1}-\sigma_{n}$ as the singular value spacings. Consequently, the singular-value-spacing ratio $r_{\upsigma,n}$ is defined analogous to the $r$-ratio as Kawabata et al. (2023)
$r_{\upsigma,n}=\frac{\mathrm{min}(\lambda_{n},\lambda_{n+1})}{\mathrm{max}(\lambda_{n},\lambda_{n+1})}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \braket{r_{\upsigma}}=\mathrm{mean}(r_{\upsigma,n})\,.$ (3)
Referred to as the $\braket{r_{\upsigma}}$-value, they adhere to statistical distributions akin to those observed for non-Hermitian Hamiltonians, similar to the $\braket{r}$-value distributions for Hermitian matrices. This framework has been instrumental in addressing the $38$-fold symmetry classification in non-Hermitian systems Kawabata et al. (2023).
$\mathrm{System}$ $N=20$ $N=22$ $N=24$ $N=26$ $N=28$ $N=30$
$\langle r\rangle_{\mathrm{RMT}}$ 0.6744 0.5996 0.5307 0.5996 0.6744 0.5996
$\langle r_{\upsigma}\rangle_{\mathrm{nSYK}}$ 0.6744 0.5997 0.5307 0.5996 0.6745 0.5995
Table 1: The $\braket{r_{\upsigma}}$-values for the non-Hermitian dense SYK model for different system sizes $N=20\,(10000)$, $N=22\,(6000)$, $N=24\,(2000)$, $N=26\,(1000)$, $N=28\,(100)$, and $N=30\,(50)$, where the parenthesis includes the number of samples taken. Based on the $\braket{r_{\upsigma}}$ values, we identify the periodicity as $N$ mod 8 = 0 (GOE), 2, 6 (GUE), and 4 (GSE) You et al. (2017). These GOE, GUE, and GSE correspond to the eigenvalue statistics of the Hermitian matrices.
The values are in precise agreement (within numerical accuracy) with the $\braket{r}_{\mathrm{RMT}}$-values for large $N$ results in Atas et al. (2013).
Table  1 showcases the $\braket{r_{\upsigma}}$-value for the dense non-Hermitian SYK model for varying system sizes $N$, representing the total number of fermions. The data is juxtaposed with corresponding eigenvalue statistics of Hermitian Gaussian ensembles, revealing a striking correlation emphasizing the utility of the $\braket{r_{\upsigma}}$-value in non-Hermitian systems. Similar to its Hermitian counterpart, we observe an $N$ mod 8 periodicity using singular value statistics: $N$ mod 8 = 0 corresponds to GOE, $N$ mod 8 = 2 and 6 corresponds to GUE, and $N$ mod 8 = 4 corresponds to GSE for the eigenvalue statistics of the Hermitian matrices. Figure 1 correspondingly shows the consecutive singular level-spacing distribution $p_{\sigma}(\lambda)$ for $N=24$ (GOE), $N=26$ (GUE) and $N=28$ (GSE) respectively, which are approximated by the analogous Hermitian Wigner-Dyson distribution Mehta (1991). Note that we are comparing the singular value statistics of non-Hermitian matrices to the eigenvalue statistics of Hermitian matrices. The singular value correlations in Hermitian matrices are generally weaker than the eigenvalue correlations and exhibit different spacing ratios Kawabata et al. (2023). However, the nSYK falls into the different non-Hermitian symmetry classes, which can also be used to show the validity of the singular value statistics García-García et al. (2022a).
Figure 1: The consecutive singular level-spacing distribution in the non-Hermitian dense SYK model for $N=24\,(2000)$ (GOE), $26\,(1000)$ (GUE), and $28\,(100)$ (GSE) with the corresponding realizations shown in the brackets. The solid lines represent distribution for GOE, GUE, and GSE respectively taking matrices of size $10^{4}\times 10^{4}$ with averaging over 1000 realizations.
SVD in sparse non-Hermitian SYK model: We now turn our attention to the sparse variant of the model, which is governed by the sparsity parameter $p$. Figure 2 illustrates the dependence of $p$ on $\braket{r_{\upsigma}}$-value for different system sizes. Each system size, representing a distinct ensemble, is depicted in a unique color. The dotted lines mark the expected values for Poisson statistics and the Hermitian ensemble, as specified in Table 1. It is observed that the Hamiltonian (1) retains its chaotic nature up to a critical level of sparsity denoted by $p_{\mathrm{crit}}$. This chaotic behavior is also corroborated by the $\braket{r_{\upsigma}}$-ratio presented in Table 1. Consequently, we define a transition point $p_{\mathrm{crit}}$ beyond which the Hamiltonian (1) ceases to exhibit chaotic dynamics. Such transition point is obtained by considering approximately $99\%$ of the corresponding $\braket{r_{\upsigma}}$ value at $p=1$ Orman et al. (2024). Notably, as the system size increases, this transition point migrates towards lower sparsity levels. When the sparsity is sufficiently low, the $\braket{r_{\upsigma}}$-value aligns with the Poisson limit, signifying the Hamiltonian’s transition to complete integrability.Note that when a term of the Hamiltonian very weakly breaks a degeneracy, the resulting spectrum has $r_{\sigma}$ values close to zero, which can cause the $\braket{r_{\upsigma}}$ value to fall below the Poisson limit.
In conjunction with the above, the relationship between the critical sparsity $p_{\mathrm{crit}}$ and the system size $N$ is also noteworthy. The critical value diminishes as the system size increases. This reduction follows a power-law trend, best described by fitting the data to the curve Orman et al. (2024)
$p_{\mathrm{crit}}=kN\binom{N}{4}^{-1}\approx\frac{24k}{N^{3}}\,,$ (4)
in the large $N$ limit. where the decrease is proportional to $1/N^{3}$ and the constant $k$ is approximately $1.68$, as determined by the fitting process. This scaling behavior is consistent with that observed for the sparse Hermitian SYK model, as reported in García-García et al. (2021); Orman et al. (2024).
Figure 2: (Top) Variation of $\braket{r_{\upsigma}}$-value with various level of sparsity $p$. The size of the circle varies with $N$ i.e., the largest $N$ corresponds to the leftmost plot for each color. At low sparsity, the $\braket{r_{\upsigma}}$ matches with the corresponding random matrix ensemble until a threshold $p_{\mathrm{crit}}$ is reached, after which $\braket{r_{\upsigma}}$ drops to the Poissonian value, indicating a hint of integrability. The estimated errors for the values obtained are smaller than the markers in the plot. (Bottom) The finite-size dependence of the threshold or the critical value $p_{\mathrm{crit}}$. It decreases as the system size increases. The decrease scales as $1/N^{3}$ with large system size, as given in (4).
Spectral and Singular form factor: Contrary to the singular value statistics, the SFF probes the long-range eigenvalue correlations. It is defined as the Fourier transform of two-point eigenvalue correlation pairs Leviandier et al. (1986); Wilkie and Brumer (1991); Brézin and Hikami (1997). For systems exhibiting quantum chaos, the SFF displays a characteristic dip-ramp-plateau pattern. The initial dip, evident in the early-time regime, is a non-universal trait influenced by the detailed nature of the energy spectrum Agam et al. (1995). This dip extends until a timescale known as the Thouless time ($t_{\mathrm{Th}}$) Gharibyan et al. (2018); Nosaka et al. (2018) or ramp time Cotler et al. (2017) beyond which the energy difference surpasses the inverse of $t_{\mathrm{Th}}$, referred to as the Thouless energy Thouless (1977). Subsequently, a ramp emerges at $t_{\mathrm{Th}}$, indicative of eigenvalue repulsion, a phenomenon that mirrors the SFF behavior predicted by RMT Brézin and Hikami (1997); Cotler et al. (2017) and underscores the chaotic dynamics of the Hamiltonian. Moreover, this ramp hints at the presence of a holographic dual to the corresponding theory. Beyond the Heisenberg time Erdős and Knowles (2015); Prange (1997), which scales inversely to the mean level spacing, the SFF reaches a plateau that depends on the system size. In contrast, integrable systems typically bypass the ramp stage and saturate directly from the dip to the plateau regime. SFF in sparse SYK models has been thoroughly investigated in recent literature García-García et al. (2021); Tezuka et al. (2023); Anegawa et al. (2023); Orman et al. (2024). Several other generalizations for SFF in non-Hermitian and open systems have also been proposed García-García et al. (2023a); Xu et al. (2021); Zhou et al. (2024).
Analogous to the SFF, the singular form factor ($\upsigma$FF) is defined as Roccati et al. (2024)
$\upsigma\mathrm{FF}(t)=\frac{1}{L^{2}}\braket{}{Z_{\upsigma}(it)}{{}^{2}}=\frac{1}{L^{2}}\bigg{\langle}\Big{|}\sum_{n}e^{-i\sigma_{n}t}\Big{|}^{2}\bigg{\rangle}\,,$ (5)
where $Z_{\upsigma}(it)=\sum_{n}e^{-i\sigma_{n}t}$ is the infinite-temperature partition function of singular values, which can be further expressed as the overlap between the time-evolved right or left singular vectors governed by the effective Hamiltonian $\sqrt{H^{\dagger}H}$ or $\sqrt{HH^{\dagger}}$ respectively Roccati et al. (2024). Here the angle brackets indicate an ensemble average, which is particularly relevant in the context of a disordered system since $\upsigma$FF is not self-averaging Prange (1997).
Figure 3: Behavior of $\upsigma$FF$(t)$ for various sparsity parameter $p$. The linear ramp ceases to exist as the sparsity increases. The system parameters are $N=26$ (GUE) with $1000$ Hamiltonian realizations. Here we choose the Gaussian filter with $\alpha=3.27$, which lies within the prescribed range in Orman et al. (2024).
Figure 3 shows the distinct dip-ramp-plateau pattern exhibited by $\upsigma$FF for the Hamiltonian (1), across varying degrees of sparsity. In this work, we consider a system size of $N=26$ and analyze the outcomes from $1000$ independent realizations of the Hamiltonian to ensure robust statistical analysis. To reduce the oscillations in the dip region, we use the filter function $Y_{\upsigma}(\alpha,t)$ Gharibyan et al. (2018) such that
$\upsigma\mathrm{FF}(t)=\bigg{\langle}\frac{|Y_{\upsigma}(\alpha,t)|^{2}}{|Y_{\upsigma}(\alpha,0)|^{2}}\bigg{\rangle}\,,$ (6)
Here the Gaussian filter function for the singular values is defined as $|Y_{\upsigma}(\alpha,t)|^{2}=|\sum_{n}e^{-\alpha\sigma_{n}^{2}}\,e^{-i\sigma_{n}t}|^{2}$. Similar to the eigenspectrum Gharibyan et al. (2018); Nosaka et al. (2018), it ensures that the dominant contribution is received from the bulk of the singular value spectrum. Here $\alpha$ is an $N$-dependent parameter that controls the Gaussianity of the spectrum. This filtering is performed for the usual spectrum, without unfolding it. For the latter case, an equivalent of connected unfolded SFF García-García et al. (2018a); Nosaka et al. (2018) can be defined. We employ the filter function on the singular values by choosing $\alpha=3.27$, which lies inside the range given in Orman et al. (2024) for $N=26$. Filter functions for SFF have also recently been shown to be associated with quantum channels Matsoukas-Roubeas et al. (2023b).
A notable observation in Fig. 3 is a parallelism between the behavior of $\upsigma$FF in non-Hermitian systems and the SFF in Hermitian systems. This similarity underscores the efficacy of $\upsigma$FF as a robust indicator for quantum chaos in non-Hermitian frameworks. In regimes of lower sparsity, the presence of the ramp is evident, signifying the retention of chaotic dynamics within this parameter space. It remains pronounced until a critical sparsity threshold is reached. Beyond this juncture, the linear ramp—a signature trait of non-integrable systems—ceases to manifest, indicating a shift towards integrability and a departure from chaotic behavior.
Figure 4 shows the variation of Thouless time ($t_{\mathrm{Th}}$), for $N=26$, with sparseness parameter $p$. The $t_{\mathrm{Th}}$ is calculated by considering the time at which the $\upsigma\mathrm{FF}$ of the system considered first intersects the SFF of the GUE random matrices Cotler et al. (2017). To determine the intersection point precisely, we consider the fractional error function $\epsilon(t):=\Big{|}\frac{\upsigma\mathrm{FF}(t)-\upsigma_{\mathrm{ramp}}(t)}{\upsigma_{\mathrm{ramp}}(t)}\Big{|}$ Gharibyan et al. (2018) indicating the deviation of the computed $\upsigma\mathrm{FF}$ to the linear-fitted function for the ramp $\upsigma_{\mathrm{ramp}}(t)$ for $p=1$. We consider the value $t$ as the Thouless time or ramp time $t_{\mathrm{Th}}$, where the fractional error $\epsilon(t)$ reaches $20\%$. The red line shows the fitting $t_{\mathrm{Th}}\approx a/p^{b}+c$ with $a=1.923$, $b=0.8767$ and $c=0.007$. Curiously, this fitting follows closely to the $1/p$ scaling in contrast to the $1/p^{2}$ scaling in Hermitian systems Orman et al. (2024). Such distinction is possible since the singular values and the eigenvalues are inherently different (see the Supplemental Material sup S1). Yet, we do not rule out the possibility of the scaling of the exponent $b$ with $N$.
Figure 4: (Main plot) Thouless time (ramp time) with respect to sparsity parameter. The red line is the fit $t_{\mathrm{Th}}\approx a/p^{b}+c$ with $a=1.923$, $b=0.8767$ and $c=0.007$. (Inset) The behavior of SFF and the $\upsigma$FF for the Hermitian dense SYK4 model. The system parameters are $N=26$ with $1000$ random Hamiltonian realizations taken. The $\upsigma$FF (green line) is computed considering all the positive singular values. It reduces to the SFF after taking the appropriate rescaling of the singular vectors and singular values, which otherwise equals the eigenvalues (with appropriate signs).
Spectral complexity for singular values - Singular complexity: Within the framework of the AdS/CFT correspondence Maldacena (1998); Witten (1998), a particular spectral measure has been proposed to capture the essence of long-range spectral correlations. This measure, known as spectral complexity, is proposed to be a dual quantity of the Einstein-Rosen bridge Iliesiu et al. (2022) in the gravitational theory, serving as an analog for the computational and circuit complexity Jefferson and Myers (2017). In a similar spirit, we define singular complexity at infinite temperature as
$C_{\upsigma}(t)=\frac{1}{L^{2}}\sum_{\sigma_{i}\neq\sigma_{j}}\left[\frac{\sin(t(\sigma_{i}-\sigma_{j})/2)}{(\sigma_{i}-\sigma_{j})/2}\right]^{2}\,,$ (7)
designed to be suitable for non-Hermitian systems. At early times, it grows quadratically, followed by a linear growth and plateau regime (see Supplemental Material sup S2 for more details). The early time behavior mirrors the behavior of spread complexity Balasubramanian et al. (2022); Erdmenger et al. (2023) based on the Krylov space approach Parker et al. (2019). The plateau value can be obtained analytically by taking the long-time average
[…]
which only depends on the difference between the singular values. In the Hermitian case, such saturation values have been studied for quantum billiards Camargo et al. (2024a) and mixed-field Ising model Camargo et al. (2024b), capturing the spectral rigidity. Integrable systems saturate at higher values compared to chaotic ones due to the presence of level repulsion, as evident from the equivalent eigenspectrum of (8).
Figure 5: Behavior of the saturation of the singular complexity with different sparseness for $N=26$ with $1000$ random Hamiltonian realizations.
Figure 5 depicts the saturation of singular complexity at varying levels of sparsity, which maintains a magnitude order of $O(10^{4})$ until it encounters a critical threshold. Beyond this threshold, there is a marked escalation at lower sparsity levels signaling a level-repulsion between the singular values. This critical value is dependent on the system size $N$, and for $N=26$, it occurs within the sparsity regime $0.003\lesssim p_{\mathrm{crit}}\lesssim 0.004$. This value resembles the expected $\braket{r_{\upsigma}}$-value shown in Fig. 2.
Exhibiting chaotic dynamics is considered to be a crucial constraint that possesses a holographic dual geometry. While this discussion does not explore the specifics of such geometry, some proposals exist in the literature García-García and Godet (2021); García-García et al. (2023b); Cai et al. (2022); Cao and Ge (2024). The breakdown of chaotic nature suggests that such dual geometry corresponding to a non-Hermitian system may be compromised at low sparsity.
Conclusion and Outlook: Drawing on the foundational work Kawabata et al. (2023) on the applicability of SVD in non-Hermitian systems, our study addresses the quantum chaotic attributes of the sparse SYK model across varying levels of sparsity. The complex eigenvalues inherent to non-Hermitian Hamiltonians pose challenges to the generalization of traditional eigenvalue statistics. This complexity persists despite strides made with complex spacing ratios Sá et al. (2020) and the DSFF Li et al. (2021, 2024). Nonetheless, recent advancements propose a novel perspective by focusing on singular values as opposed to eigenvalues within the non-Hermitian framework. Unlike their complex counterparts, singular values are inherently real and non-negative, aligning with the real eigenvalues in Hermitian systems upon appropriate alignment with singular vectors, as detailed in the Supplemental Material sup S1.
Our investigation underscores that the statistical analysis of singular values, which elucidates short-range correlations, coupled with the $\upsigma$FF and singular complexity, illuminating long-range correlations among singular values, serve as potent indicators of the chaos within non-Hermitian systems. As such, exploring the holographic bulk dual of singular complexity may offer insights into a potential holographic counterpart for non-Hermitian systems, especially the effective range of $k$ in (4) in relation to $2d$ de Sitter (dS2) gravity García-García et al. (2023b), which we leave for future work. Moreover, these tools adeptly reflect the shift towards integrability as sparsity intensifies. Such transition differs from other chaotic-integrable transitions by two-point interactions observed in the SYK model García-García et al. (2018b). We envision that these features will be exhibited in other correlation measures such as higher spacing ratios Srivastava et al. (2018); Tekur et al. (2018); Shir et al. (2023) and the number variance García-García et al. (2023a); Gharibyan et al. (2018); Evers and Mirlin (2008); García-García and Verbaarschot (2016).
Parallel research avenues have explored the efficacy of sparse Hamiltonians Xu et al. (2020); Tezuka et al. (2023) within quantum processors Jafferis et al. (2022); Kobrin et al. (2023), maintaining chaotic dynamics to a substantial degree of sparsity. Notably, the SYK model retains its holographic duality even at minimal sparsity levels García-García et al. (2021); Orman et al. (2024). Our study contributes to the burgeoning field of non-Hermitian systems in the realm of open-system dynamics, proposing a framework for the extrapolation of quantum chaotic properties within these systems. This might serve as the benchmark results for the simulation of sparse non-Hermitian systems on quantum processors.
Acknowledgements: We would like to thank Hugo A. Camargo, Adolfo del Campo, Aurélia Chenu, Antonio M. García-García, Luca V. Iliesiu, Yiyang Jia, Hosho Katsura, and Tadashi Takayanagi for fruitful discussions, comments, and suggestions on the draft. Numerical computations were performed in the Sushiki workstation using the computational facilities of YITP. P.N. thanks the Berkeley Center for Theoretical Physics (BCTP), University of California for hosting him through the Adopting Sustainable Partnerships for Innovative Research Ecosystem (ASPIRE) program of Japan Science and Technology Agency (JST), Grant No. JPMJAP2318 during the final stages of the work. This work is supported by the Japan Society for the Promotion of Science (JSPS) Grants-in-Aid for Transformative Research Areas (A) “Extreme Universe” No. JP21H05190 (P.N.) and JP21H05182, JP21H05185 (M.T.). The Yukawa Research Fellowship of T.P. is supported by the Yukawa Memorial Foundation and JST CREST (Grant No. JPMJCR19T2). The work of M.T. was partially supported by the JSPS Grants-in-Aid for Scientific Research (KAKENHI) Grants No. JP20K03787 and JST CREST (Grant No. JPMJCR24I2).
Authors’ names are listed in alphabetical order.
Supplemental Materials:
Probing quantum chaos through singular-value correlations
in sparse non-Hermitian SYK model
Pratik Nandy, Tanay Pathak, and Masaki Tezuka
Appendix A S1. From $\upsigma$FF to SFF: Reduction to Hermitian systems
In Hermitian systems, a key observation is that the singular values correspond to the absolute values of the eigenvalues. This distinction is crucial, as it implies that the SFF and $\upsigma$FF, exhibit fundamentally different behaviors. This difference is illustrated in the inset of Fig. 4, which depicts the dynamics of SFF and $\upsigma$FF for the dense Hermitian SYK4 model. The model is derived by setting $x_{abcd}=1$ and $M_{abcd}=0$ for all $\{a,b,c,d\}$ in its non-Hermitian counterpart (Eq. (1)). Replacing $\sigma_{n}$ by eigenvalues $E_{n}$ in Eq. (6), we get the analogous definition of filtered SFF.
      $\mathrm{System}$   $\braket{r_{\upsigma}}$   $\braket{r}$  $\braket{r_{+}}$  $\braket{r_{-}}$ $\braket{r}$ (2 blocks)
nSYK ($N=20$) 0.4119 0.6743 0.6744 0.6744 0.4117
nSYK ($N=22$) 0.4220 0.5994 0.5998 0.5991 0.4220
nSYK ($N=24$) 0.4238 0.5303 0.5303 0.5302 0.4234
Table 2: The Table shows the comparison between the singular value $\braket{r_{\upsigma}}$-value and the $\braket{r}$-values for the dense Hermitian SYK model for different system sizes $N=20$ (GSE), $N=22$ (GUE), $N=24$ (GOE). We take 1000 Hamiltonian realizations for each case. The Hermitian model is obtained by setting $x_{abcd}=1$ and $M_{abcd}=0$ for all $\{a,b,c,d\}$ in the Hamiltonian (1). Here the $\braket{r_{+}}$-value includes the positive eigenvalues while $\braket{r_{-}}$-value includes the negative eigenvalues only. They are to be matched with the $\braket{r}_{\mathrm{RMT}}$ values in Table 1 and Atas et al. (2013). The $\braket{r_{\upsigma}}$-values are compared with the $\braket{r}$-value of $2$ blocks (marked in blue) Orman et al. (2024). Explanation is given in the text.
Notably, the singular values offer a weaker effect compared to the eigenvalues, considering the ramp. However, to reduce the SFF from $\upsigma$FF, we consider the following prescription. From SVD, we obtain two sets of vectors, the left and right singular vectors. They are related by
[…]
where $\{\sigma_{n}\}$ are the singular values. For Hermitian systems, the eigenvalues are real, and the left (|$\ket{u_{n}}$) and the right (|$\ket{v_{n}}$) singular vectors are related with a factor of ±$\pm 1$. If the left and right singular vectors are equal, i.e., |$\ket{u_{n}}=\ket{v_{n}}$ for some $n$, then the corresponding singular value exactly equals the eigenvalue i.e., $\sigma_{n}=E_{n}$. However, if they differ by a negative sign i.e., |$\ket{u_{n}}=-\ket{v_{n}}$ for some $n$, then the corresponding singular value and the eigenvalue differs by a negative sign i.e., $\sigma_{n}=-E_{n}$. It is thus straightforward to identify such vectors such that the singular values can be directly mapped with the eigenvalues, and correspondingly compute the $\upsigma$FF. In this case, the $\upsigma$FF exactly equals the SFF as shown by the red line in the inset of Fig. 4.
Intriguingly, the Hermitian system exhibits a unique characteristic in its $\braket{r_{\upsigma}}$-value. As depicted in Table 2, there is a notable distinction when comparing the $\braket{r_{\upsigma}}$-value to the $\braket{r}$-values within the identical Hermitian SYK4 model. Additionally, the $\braket{r_{+}}$-value and $\braket{r_{-}}$-value were computed separately, considering only the positive and negative eigenvalues of the Hamiltonian, respectively. These ratios align with the overall $\braket{r}$-value, which is anticipated since it solely accounts for the spacing between consecutive eigenvalues. Nonetheless, these individual ratios as well as the overall $\braket{r}$-value do not correspond to the $\braket{r_{\upsigma}}$-value, which is identified as the $\braket{r}$-value for two separate blocks before symmetry resolution. This discrepancy arises because, in Hermitian systems, the singular values are determined by the absolute magnitudes of the eigenvalues, leading to the computation of the $\braket{r_{\upsigma}}$-value mirroring the $\braket{r}$-value for bifurcated blocks Giraud et al. (2022); Kawabata et al. (2023); Orman et al. (2024). We anticipate such bifurcation has a similar effect of having the presence of conserved charges Iyoda et al. (2018); Sá and García-García (2022). Curiously, this phenomenon is exclusive to Hermitian and anti-Hermitian (with $J_{abcd}=0$, $M_{abcd}\neq 0$) systems and echoes the discrepancy between the $\upsigma$FF and SFF as shown in the inset of Fig. 4.
Appendix B S2. Behavior of singular complexity
Figure 6: Behaviour of the singular complexity with time for dense ($p=1$) non-Hermitian SYK model. The system parameter is $N=26$ with $1000$ random Hamiltonian realizations taken. Inset shows the early time behavior of the complexity, which increases quadratically followed by linear growth and saturation.
Here we present the detailed behavior of singular complexity Iliesiu et al. (2022). It is defined at finite temperature $T=1/\beta$ as
$C_{\upsigma}(t)=\frac{1}{Z_{2\upsigma}(\beta)L}\sum_{\sigma_{i}\neq\sigma_{j}}\left[\frac{\sin(t(\sigma_{i}-\sigma_{j})/2)}{(\sigma_{i}-\sigma_{j})/2}\right]^{2}e^{-\beta(\sigma_{i}+\sigma_{j})}\,.$ (10)
Here $Z_{\upsigma}(\beta)=\sum_{n}e^{-\beta\sigma_{n}}$ is the thermal singular partition function, constructed from the singular values of the corresponding non-Hermitian Hamiltonian. In the case of degeneracies, we keep the degenerate spectrum. The derivative of the singular complexity is related to the $\upsigma$FF as
$\frac{d^{2}}{dt^{2}}C_{\upsigma}(t)=\frac{2}{L}\frac{Z_{\upsigma}(\beta)^{2}}{Z_{\upsigma}(2\beta)}\upsigma\mathrm{FF}(t)-\frac{2}{L}\,,$ (11)
in a similar spirit to spectral complexity Iliesiu et al. (2022); Erdmenger et al. (2023) which can be checked. Focusing on the infinite temperature $\beta=0$ limit, (10) takes a simpler form as in the main text:
$C_{\upsigma}(t)=\frac{1}{L^{2}}\sum_{\sigma_{i}\neq\sigma_{j}}\left[\frac{\sin(t(\sigma_{i}-\sigma_{j})/2)}{(\sigma_{i}-\sigma_{j})/2}\right]^{2}\,.$ (7)
At early times, it grows quadratically as $C_{\upsigma}(t)\approx(1-1/L)t^{2}$, followed by a linear growth and plateau regime. Figure 6 shows the time evolution of the infinite-temperature singular complexity (7) for the dense Hamiltonian for $N=26$ and $1000$ realizations. The complexity grows quadratically, followed by linear growth (inset) and saturation. The saturation value is elevated as the sparsity increases.